### v1

# Joint Velocity-Growth Flow Matching for Single-Cell Dynamics Modeling

Abstract
Learning the underlying dynamics of single cells from snapshot data has gained increasing attention in scientific and machine learning research. The destructive measurement technique and cell proliferation/death result in unpaired and unbalanced data between snapshots, making the learning of the underlying dynamics challenging. In this paper, we propose joint Velocity-Growth Flow Matching (VGFM), a novel paradigm that jointly learns state transition and mass growth of single-cell populations via flow matching. VGFM builds an ideal single-cell dynamics containing velocity of state and growth of mass, driven by a presented two-period dynamic understanding of the static semi-relaxed optimal transport, a mathematical tool that seeks the coupling between unpaired and unbalanced data. To enable practical usage, we approximate the ideal dynamics using neural networks, forming our joint velocity and growth matching framework. A distribution fitting loss is also employed in VGFM to further improve the fitting performance for snapshot data. Extensive experimental results on both synthetic and real datasets demonstrate that VGFM can capture the underlying biological dynamics accounting for mass and state variations over time, outperforming existing approaches for single-cell dynamics modeling.
1 Introduction
Inferring the latent dynamics of complex systems from sparse and noisy data is a fundamental challenge in science and engineering. In many domains, e.g., stock markets [1], climate systems [2], and biological processes [3, 4, 5, 6, 7, 8, 9, 10], continuous trajectories are rarely fully observed by sensors. Instead, cross-sectional snapshot data collected at discrete time points are commonly provided. This challenge is especially important in single-cell RNA sequencing [11, 12, 13, 14, 15], where destructive sampling yields unpaired population-level snapshots across time without having the tracked individual cell fates. Furthermore, due to the mass changes during cellular development or response processes, observed data often exhibits mass unbalancedness across time, violating mass conservation. Consequently, reconstructing the time-evolving, unnormalized density function from limited samples has become an important research problem and attracted increasing attention [13, 14, 16].
Deep learning-based models for dynamics inference have demonstrated great potential [17, 3, 18, 7, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28]. These models typically employ ordinary or stochastic differential equations (ODEs or SDEs) parameterized by neural networks to approximate the velocity field governing density evolution. One class of methods, based on simulation [29, 17, 3, 30, 23, 19, 18, 22, 25], generates synthetic trajectories by feeding initial data through neural networks and solving ODEs or SDEs numerically, and compares simulated results with observations for loss computation. However, these simulation-based methods rely heavily on numerical solvers during training, which significantly increases computational cost. In high-dimensional settings, the enlarged search space further exacerbates instability, hindering scalability and convergence. Another line of research focuses on simulation-free approaches [31, 32, 33, 34, 21, 35, 36, 37, 38, 39, 40, 28, 41], where the velocity field is trained efficiently by constructing conditional probability paths without trajectory simulation, yielding better efficiency and stability in training compared with simulation approaches.
Nevertheless, the simulation-free approaches mentioned above rely solely on the velocity field $v$, ignoring the unbalancedness of observed data, which can result in incorrect reconstruction of underlying dynamics or unsatisfying generation performance [25, 35, 40]. Distributional imbalance is a common phenomenon in single-cell dynamics, where cellular proliferation and death occur, thus necessitating the introduction of a growth term $g$ to allow for mass increasing or decreasing. To address this, [42, 43] proposed to jointly learn velocity $v$ and growth $g$ by minimizing the Wasserstein-Fisher-Rao (WFR) metric [44, 45]. However, it mathematically enforces $v=\nabla g$, which lacks clear mechanistic justification in biological systems. Other WFR-inspired approaches [22, 25] design separate neural networks for $v$ and $g$ and prioritize fitting the distributions. However, these methods require heavy simulation during training, limiting their scalability to high dimensions.
Figure 1: The goal of this paper is to learn a joint state transition (controlled by $v_{\theta}(\mathbf{x},t)$) and mass growth (controlled by $g_{\omega}(\mathbf{x},t)$) dynamics for single-cell evolution.
In this paper, we propose joint Velocity-Growth Flow Matching (VGFM), a novel approach for single-cell dynamics modeling. VGFM aims to learn the joint state transition [^1] State transition describes changes in gene expression in a cell. and mass growth [^2] Mass growth means mass increasing or decreasing. of single-cell evolution by flow matching, as illustrated in Fig. 1. VGFM is based on the static semi-relaxed optimal transport that allows mass variation in building coupling between snapshots, for which we present a dynamic understanding that enables mass growth and state transition accomplished in two time periods, respectively. Based on this dynamic understanding, we build a joint transition and growth dynamics, inheriting the optimal properties of semi-relaxed optimal transport. We then approximate the velocity and growth in the built dynamics using neural networks with finite samples, yielding a joint velocity and growth flow matching framework. To further improve fitting performance for the single-cell snapshots, we employ a distribution fitting loss based on the Wasserstein distance. Thanks to the joint velocity and growth flow matching, our approach enjoys more stable training and better scalability to high dimensions, compared with simulation-based approaches. Extensive experiments on both synthetic and real single-cell datasets are conducted to evaluate our proposed approach. Experimental results demonstrate accurate dynamics reconstruction and superior performance achieved by VGFM over recent baselines.
2 Background and Related Works
This section revisits optimal transport, and reviews the flow-matching-based methods for single-cell and unbalancedness-aware distribution learning methods, which are mostly related to this paper.
Optimal transport.
Given two distributions $p_{0}$ and $p_{1}$ on domain $\Omega$ satisfying $\int_{\Omega}p_{0}(\mathbf{x})\,\mathrm{d}x=\int_{\Omega}p_{1}(\mathbf{x})\,\mathrm{d}x$, the optimal transport [46] aims to find a joint distribution of $p_{0}$ and $p_{1}$, named coupling, such that the transport cost is minimized, formulated as the following optimization problem:
$\min_{\pi\in U(p_{0},p_{1})}\int_{\Omega^{2}}c(\mathbf{x}_{0},\mathbf{x}_{1})\,\mathrm{d}\pi(\mathbf{x}_{0},\mathbf{x}_{1}),\mbox{ s.t.}\ U(p_{0},p_{1})=\left\{\pi\geq 0:{\rm P}_{\#}^{0}\pi=p_{0},\ {\rm P}_{\#}^{1}\pi=p_{1}\right\},$ (1)
where $c$ is the cost function. For a map $T$, we define $T_{\#}p_{0}(\mathbf{x})=\int_{\mathbf{x}^{\prime}:T(\mathbf{x}^{\prime})=\mathbf{x}}p_{0}(\mathbf{x}^{\prime})\mathop{}\!\mathrm{d}\mathbf{x}^{\prime}$. Let ${\mbox{P}}^{i}(\mathbf{x}_{0},\mathbf{x}_{1})$ $=\mathbf{x}_{i}$ for $i=0,1$, then we have ${\rm P}_{\#}^{i}\pi(\mathbf{x}_{i})=\int_{\Omega}\pi(\mathbf{x}_{0},\mathbf{x}_{1})\,\mathrm{d}\mathbf{x}_{i}$ for $i=0,1$. When $c(\mathbf{x}_{0},\mathbf{x}_{1})=\|\mathbf{x}_{0}-\mathbf{x}_{1}\|^{2}$, by Brenier’s theorem [47], the optimal coupling can be expressed as $\pi^{*}=(\mathrm{Id},T^{*})_{\#}p_{0}$, where $T^{*}$ is called the Monge map.
Flow matching for learning single cell dynamics.
Flow matching [31, 32, 33] is a simulation-free approach where the velocity field is trained efficiently by constructing conditional probability paths without trajectory simulation. Building on this foundation, further improvements have been achieved by incorporating optimal transport guidance [34, 21], considering unbalancedness [35], accounting for manifold structures [36], and approximating Schrödinger bridge via flow and score matching [39], all of which has been applied to model single-cell dynamics [21, 39, 36, 35]. Different from these methods, our method explicitly models the simultaneous matching of both the velocity field and the growth function, driven from our developed ideal joint state transition and mass growth dynamics, which does not require the construction of conditional probability paths like the above approaches.
Unbalancedness-aware distribution learning methods.
Data imbalance is prevalent across a variety of domains, such as image synthesis and protein generation, motivating the development of flow matching models that account for unbalanced distributions [35, 40, 48]. However, these methods often do not model growth functions, which are essential in single-cell contexts. In the single-cell domain, several approaches have been proposed to model growth dynamics [42, 43, 5, 22, 49], yet they either fail to leverage the informative variations in cell abundance observable from snapshot data [42, 43], or rely heavily on computationally expensive simulations [5, 22, 49], limiting their scalability and efficiency. Differently, our method explicitly learns the growth function from observed snapshot data, yielding better performance for single cells than these methods as shown in experiments.
3 Method
Given the snapshot population of single cells, this paper aims to build the dynamic trajectory of single cells. Considering the unbalancedness between snapshot populations due to the undergo cell proliferation or death, we aim to build a model that can transform between unbalanced distributions by jointly learning a velocity field for controlling state transition and a growth function for controlling mass variation. Towards this goal, we propose joint Velocity-Growth Flow Matching (VGFM) based on semi-relaxed optimal transport to learn the dynamics of single cells. As illustrated in Fig. 2, we first present a decoupled understanding of state transition and mass growth for unbalanced dynamics based on semi-relaxed optimal transport (Figs. 2 (b), (d)), upon which we build a dynamic process between unbalanced distributions (Fig. 2 (c) ). Finally, we propose the velocity and growth flow matching using the built dynamic process to learn the velocity and growth with samples (Fig. 2 (a)). Next, we first describe the unbalanced dynamics of single-cell data, and then discuss the details of each component of our method.
3.1 Unbalanced Dynamics of Single Cell
Given two adjacent snapshots of populations/distributions denoted as $p_{0}$ and $p_{1}$, if the system is balanced, the dynamics $\bar{p}_{t}(t\in[0,1])$ between $p_{0}$ and $p_{1}$ can be generated by ODE $\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}=v_{t}(\mathbf{x}_{t})$ such that $\mathbf{x}_{0}\sim p_{0}$ and $\mathbf{x}_{1}\sim p_{1}$. Corresponding to this ODE, $\bar{p}_{t}$ is governed by continuity equation $\partial_{t}\bar{p}_{t}=-\nabla\cdot(\bar{p}_{t}v_{t})$. However, in biological systems, cells can proliferate and die. Such behaviors can be captured by a time-dependent weight $w_{t}(\mathbf{x}_{t})$ associated with the state transition, whose evolution is controlled by a growth function $g$, simulating cell proliferation or death processes:
$\begin{cases}&\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}=v_{t}(\mathbf{x}_{t}),\\ &\frac{\mathrm{d}\log w_{t}(\mathbf{x}_{t})}{\mathrm{d}t}=g_{t}(\mathbf{x}_{t}).\end{cases}$ (2)
Under this formulation, the unbalanced distribution dynamics $p_{t}$ is generated jointly by the position $\mathbf{x}_{t}$ and the weight $w_{t}(\mathbf{x}_{t})$. Specifically, $w_{t}(\mathbf{x}_{t})$ models the variation of $p_{t}$ against $\bar{p}_{t}$, i.e., $w_{t}(\mathbf{x}_{t})=p_{t}(\mathbf{x}_{t})/\bar{p}_{t}(\mathbf{x}_{t})$ for $\bar{p}_{t}(\mathbf{x}_{t})\neq 0$. Corresponding to Eq. (2), ${p}_{t}$ is governed by
$\partial_{t}p_{t}=-\nabla\cdot(p_{t}v_{t})+g_{t}p_{t}.$ (3)
Figure 2: Illustration of our proposed VGFM, consisting of the velocity and growth flow matching deduced by the dynamic reformulation of the semi-relaxed optimal transport.
3.2 Building Unlabalnced Dynamics Based on Semi-Relaxed Optimal Transport
Though Eqs. (2) or (3) provide a formulation of unbalanced dynamics between $p_{0}$ and $p_{1}$, the equations admit a wide range of admissible dynamics. We next seek the dynamics based on semi-relaxed optimal transport. Specifically, given the semi-relaxed optimal transport
$\min_{\pi\geq 0}\mathcal{J}_{\rm sot}(\pi)\triangleq\int_{\Omega^{2}}c(\mathbf{x}_{0},\mathbf{x}_{1})\,\mathrm{d}\pi(\mathbf{x}_{0},\mathbf{x}_{1})+\mathrm{KL}({\rm P}_{\#}^{0}\pi\|p_{0})\quad\text{subject to}\quad{\rm P}_{\#}^{1}\pi=p_{1},$ (4)
between two unbalanced distributions $p_{0}$ and $p_{1}$, we will first present a dynamic understanding of Eq. (4) that decouples the state transition and mass growth into different periods. Based on this dynamic understanding, we will build a reasonable unbalanced dynamics between $p_{0}$ and $p_{1}$, which naturally drives our velocity and growth matching to learn the dynamics with samples (see Sect. 3.3).
Dynamic understanding of semi-relaxed optimal transport.
As mentioned above, the solution of the semi-relaxed optimal transport allows both state transition and mass variation/growth. From a dynamic perspective, we understand the transport as a two-period dynamic process, in which we perform mass growth in the first period and state transition in the second period. Note that such an understanding is not practical in applications, but it provides a decoupled modeling of state transition and mass growth, which promotes a reasonable and friendly-to-matching unbalanced dynamics modeling, driving our flow matching algorithm for learning velocity and growth, as we will make clear later. More concretely, given $\lambda\in(0,1)$, we perform mass growth controlled by $g_{t}$ for $t\in[0,\lambda]$ through $\partial_{t}p_{t}=g_{t}p_{t}$ with initial and ending distributions of $p_{0}$ and $p_{\lambda}$ respectively. For $t\in(\lambda,1]$, the state transition controlled by $v_{t}$ through $\partial_{t}p_{t}=-\nabla\cdot(p_{t}v_{t})$ is performed with initial and ending distributions of $p_{\lambda}$ and $p_{1}$, respectively. We then define the following two-period transport model:
$\min_{(v_{t},g_{t})\in\mathcal{C}_{\lambda}(p_{0},p_{1})}\mathcal{J}^{\lambda}_{\rm tpt}(v_{t},g_{t})\triangleq(1-\lambda)\int_{\Omega}\int_{\lambda}^{1}p_{t}(\mathbf{x})\|v_{t}(\mathbf{x})\|^{2}\mathrm{d}t\mathrm{d}\mathbf{x}+\mathcal{H}(v_{t},g_{t},p_{t}),$ (5)
where $\mathcal{H}(v_{t},g_{t},p_{t})=\int_{\Omega}p_{0}(\mathbf{x})(e^{\int_{0}^{\lambda}g_{t}(\mathbf{x})\mathrm{d}t}(\int_{0}^{\lambda}g_{t}(\mathbf{x})\mathrm{d}t-1)+1)\mathrm{d}x$, $\mathcal{C}_{\lambda}(p_{0},p_{1})=\{(v_{t},g_{t}):\partial_{t}p_{t}=g_{t}p_{t},t\in[0,\lambda];\partial_{t}p_{t}=-\nabla\cdot(p_{t}v_{t}),t\in(\lambda,1]\}$, and $p_{0}$ and $p_{1}$ are the given distributions. The following proposition shows the relation between the semi-relaxed optimal transport model in Eq. (4) and the two-period transport model in Eq. (5).
Proposition 1.
Assume $c(\mathbf{x}_{0},\mathbf{x}_{1})=\|\mathbf{x}_{0}-\mathbf{x}_{1}\|^{2}$ and if we enforce ${\rm P}_{\#}^{0}\pi$ and $p_{0}$ to share the same support for admissible solution $\pi$ to problem (4), then we have $\min_{\pi}\mathcal{J}_{\rm sot}(\pi)=\min_{v_{t},g_{t}}\mathcal{J}^{\lambda}_{\rm tpt}$ $(v_{t},g_{t}),\forall\lambda\in(0,1)$. Moreover, for any $\lambda\in(0,1)$, given the optimal transport plan $\pi^{*}$ to problem (4), let $p_{\lambda}^{*}\triangleq{\rm P}_{\#}^{0}\pi^{*}$, then $\pi^{*}$ can be expressed as $\pi^{*}=(\mathrm{Id},T^{*})_{\#}p^{*}_{\lambda}$ where $T^{*}$ is the Monge map between $p^{*}_{\lambda}$ and $p_{1}$. Meanwhile, there exist a $g_{t}^{*}$ such that $p_{\lambda}^{*}=p_{0}(\mathbf{x})e^{\int_{0}^{\lambda}g^{*}_{t}(\mathbf{x})\mathrm{d}t}$, and a $v_{t}^{*}$ given by $v^{*}_{t}\left(\mathbf{x}+\frac{t-\lambda}{1-\lambda}(T^{*}(\mathbf{x})-\mathbf{x})\right)=\frac{T^{*}(\mathbf{x})-\mathbf{x}}{1-\lambda}$, satisfying $(v_{t}^{*},g_{t}^{*})\in\arg\min_{v_{t},g_{t}}\mathcal{J}^{\lambda}_{\rm tpt}(v_{t},g_{t})$.
The proof is provided in the Appendix A.2. Proposition 1 indicates that problems (4) and (5) share the same optimal objective function value. Meanwhile, the optimal objective function value of the two-period transport problem (5) is independent of $\lambda$. Moreover, $v_{t}^{*},g_{t}^{*}$ can be constructed using $\pi^{*}$.
Building unbalanced dynamics.
Although we have identified a scheme in which $p_{0}$ evolves into $p_{1}$ through a two-period process governed separately by $v_{t}$ and $g_{t}$ respectively, this decoupling does not align with the behavior observed in biological systems, where transport and growth typically occur simultaneously. To address this, we next build an unbalanced dynamics based on the dynamic understanding discussed above that allows velocity and growth to jointly drive the state transition and mass growth while ensuring consistency with the target distribution. Specifically, given $v_{t}$ $(t\in(\lambda,1])$, $g_{t}$ $(t\in[0,\lambda])$ of the two-period process, we define the joint velocity $\tilde{v}_{t}$ and growth $\tilde{g}_{t}$ by
$\tilde{v}_{t}(\mathbf{x})=(1-\lambda)\cdot v_{(1-\lambda)t+\lambda}(\mathbf{x}),$ (6)
$\tilde{g}_{t}(\mathbf{x})=\lambda\cdot g_{\lambda t}\left(\psi_{\tilde{v},t}^{-1}(\mathbf{x})\right),$
where $t\in[0,1]$, and $\psi_{\tilde{v},t}$ is the flow generated by $\tilde{v}$, i.e., $\psi_{\tilde{v},t}(\mathbf{x}_{0})=\mathbf{x}_{0}+\int_{0}^{t}\tilde{v}_{s}(\mathbf{x})\mathrm{d}s$.
Theorem 1.
Given the initial distribution $p_{0}$, denote the ending distribution of the two-period dynamics
$\partial_{t}p_{t}=g_{t}p_{t},t\in[0,\lambda];\quad\partial_{t}p_{t}=-\nabla\cdot(p_{t}v_{t}),t\in(\lambda,1],$ (7)
as $p_{1}$, and denote the ending distribution of the joint dynamics starting from $p_{0}$
$\partial_{t}\tilde{p}_{t}=-\nabla\cdot(\tilde{p}_{t}\tilde{v}_{t})+\tilde{g}_{t}\tilde{p}_{t},\quad t\in[0,1],\quad\tilde{p}_{0}=p_{0},$ (8)
as $\tilde{p}_{1}$, then it holds that $\tilde{p}_{1}=p_{1}$.
The proof is detailed in Appendix A.3. Theorem 1 indicates that given the same initial distribution at $t=0$, the two-period and the joint dynamics yield the same distribution at $t=1$. Since $\tilde{v}_{t},\tilde{g}_{t}$ are defined on $[0,1]$, the joint dynamics is more practical in applications, e.g., biological systems. Meanwhile, if $v_{t},g_{t}$ is defined as in Proposition 1, $\tilde{v}_{t},\tilde{g}_{t}$ will inherit the optimality properties of the semi-relaxed optimal transport, benefitting our flow matching algorithm, as shown in Sect. 3.3.
3.3 Velocity and Growth Flow Matching
[…]
$\tilde{v}_{t}(\psi_{\tilde{v},t}(\mathbf{x}_{0}))=T^{*}(\mathbf{x}_{0})-\mathbf{x}_{0},\quad\tilde{g}_{t}(\psi_{\tilde{v},t}(\mathbf{x}_{0}))=\log{\rm P}_{\#}^{0}\pi^{*}(\mathbf{x}_{0})-\log p_{0}(\mathbf{x}_{0}),\mbox{ where }\ \mathbf{x}_{0}\sim p_{0}.$ (9)
In practice, only finite samples of $p_{0}$ and $p_{1}$ can be accessed. We then aim to learn neural networks $v_{\theta}(\mathbf{x},t),g_{\omega}(\mathbf{x},t)$ to approximate $\tilde{v}_{t}(\mathbf{x}),\tilde{g}_{t}(\mathbf{x})$ by velocity and growth flow matching as
$\min_{\theta,\omega}\mathbb{E}_{\mathbf{x}_{0}}\mathbb{E}_{t}\left[\|v_{\theta}(\psi_{\tilde{v},t}(\mathbf{x}_{0}),t)-\tilde{v}_{t}(\psi_{\tilde{v},t}(\mathbf{x}_{0}))\|^{2}+\left|g_{\omega}(\psi_{\tilde{v},t}(\mathbf{x}_{0}),t)-\tilde{g}_{t}(\psi_{\tilde{v},t}(\mathbf{x}_{0}))\right|^{2}\right],$ (10)
that can generalize to new samples. We first estimate $\pi^{*}$ and $T^{*}$ using samples. Given samples $\{\mathbf{x}_{0}^{i}\}_{i=1}^{n}$ of $p_{0}$ and $\{\mathbf{x}_{0}^{i}\}_{j=1}^{m}$ of $p_{1}$, We seek the optimal transport plan $\pi^{0\to 1}\approx\pi^{*}(\mathbf{x}_{0},\mathbf{x}_{1})$ by solving the entropy-regularized semi-relaxed transport problem using Sinkhorn-based algorithm [50, 51] as
$\pi^{0\to 1}=\arg\min_{\pi\geq 0}\sum_{i,j}c_{ij}\pi_{ij}+\epsilon H(\pi)+\tau\mathrm{KL}(\pi\mathbf{1}_{m}||\mathbf{1}_{n}),\quad\text{subject to}\quad\pi^{\top}\mathbf{1}_{n}=\mathbf{1}_{m},$ (11)
where $c_{ij}=\|\mathbf{x}_{0}^{i}-\mathbf{x}_{1}^{j}\|^{2}$, $\mathbf{1}_{m}$ is the all-one vector, $H(\pi)$ denotes the negative entropy of $\pi$, and $\epsilon$ and $\tau$ are hyperparameters. With $\pi^{0\to 1}$, $T^{*}$ can be estimated using barycentric mapping $T^{*}(\mathbf{x}_{0}^{i})\approx\frac{1}{N^{i}}\sum_{j}\pi^{0\to 1}_{ij}\mathbf{x}_{1}^{j}$ where $N^{i}=\sum_{j}\pi^{0\to 1}_{ij}$. In implementation, for each $\mathbf{x}_{0}^{i}$, we sample $j$ from $(1,2,\cdots,m)$ with probability $\frac{1}{N^{i}}\pi^{0\to 1}_{ij}$, and approximate $T^{*}(\mathbf{x}_{0}^{i})\approx\mathbf{x}_{1}^{j}$. This approximation is accurate as $\epsilon\rightarrow 0$. Therefore, $\psi_{\tilde{v},t}(\mathbf{x}_{0}^{i})\approx\mathbf{x}^{i}_{0}+t(\mathbf{x}^{j}_{1}-\mathbf{x}_{0}^{i})\triangleq\mathbf{x}_{t}$ for $t\in[0,1]$, and Eq. (10) becomes
$\mathcal{L}_{\rm VGFM}(\theta,\omega)=\sum_{i=1}^{n}\sum_{j=1}^{m}\pi_{ij}^{0\to 1}\mathbb{E}_{t}\left[\left\|v_{\theta}(\mathbf{x}_{t},t)-(\mathbf{x}_{1}^{j}-\mathbf{x}_{0}^{i})\right\|^{2}+\left|g_{\omega}(\mathbf{x}_{t},t)-\log([\pi^{0\to 1}\mathbf{1}_{m}]_{i})\right|^{2}\right],$ (12)
where $[\cdot]_{i}$ is the $i$-th element. The last term in Eq. (12) is because $p_{0}(\mathbf{x}_{0}^{i})=1$, for $i=1,\cdots,n$. In experiments, $\mathcal{L}_{\rm VGFM}$ can be implemented using mini-batch samples. Specifically, we sample $(i,j)$ from $\pi^{0\to 1}$ and $t$ from $\mathcal{U}(0,1)$, then calculate the loss in square brackets. To improve robustness, we add a Gaussian noise to $\mathbf{x}_{t}$ in experiments.
3.4 Training Process
Besides the flow matching loss in Eq. (12), we also employ distribution fitting loss to improve performance further. Next, we introduce the distribution fitting loss and our training algorithm.
To achieve better performance, we incorporate the Wasserstein distance between generated samples and observed samples as part of the loss function. Specifically, let $\mathcal{X}_{t}=\{\mathbf{x}^{i}_{t}\}_{i=1}^{N_{t}}$ for $t=0,1,\cdots,T-1$ denote the observed samples at different time points, and denote $p(\mathcal{X}_{t})=\sum_{i=1}^{N_{t}}\delta_{\mathbf{x}^{i}_{t}}$. We define $\phi_{v_{\theta}}:\mathbb{R}^{d}\times\mathcal{T}\to\mathbb{R}^{d|\mathcal{T}|}$ as the trajectory mapping function parameterized by the neural network $v_{\theta}$, which takes a given starting point as the initial condition and outputs particle coordinates at time indices $\mathcal{T}$ according to ODE dynamics, where $\mathcal{T}$ denotes a set of time steps. Given an initial set $\mathcal{X}_{0}$, the model $\phi_{v_{\theta}}$ predicts particle positions at future time points in $\mathcal{T}$. Specifically, the predicted samples are computed by applying the neural ODE to $\mathcal{X}_{0}$ over the time indices $\{1,\cdots,T-1\}$, i.e., $\hat{\mathcal{X}}_{1},\cdots,\hat{\mathcal{X}}_{T-1}=\phi_{v_{\theta}}(\mathcal{X}_{0},\{1,\cdots,T-1\})$. Similarly, we define $\phi_{g_{\omega}}$ as the particle weight mapping function parameterized by the neural network $g_{\omega}$, which takes the initial weight of particle $i$ as input and outputs the corresponding weight values at $\mathcal{T}$ under ODE dynamics. For simplicity of notation, we use $\hat{w}(\hat{\mathbf{x}})\mbox{ for }\hat{\mathbf{x}}\in\hat{\mathcal{X}}_{t}$ to denote the weight generated by $\phi_{g_{\omega}}$ corresponding to the particles in set $\hat{\mathcal{X}}$ which is generated by $\phi_{v_{\theta}}$ The distribution fitting loss is defined as the Wasserstein distance as
$\mathcal{L}_{\mathrm{OT}}(\theta,\omega)=\sum_{t=1}^{T-1}\mathrm{\mathcal{W}_{1}}\left(\frac{1}{N_{t}}p(\mathcal{X}_{t}),\frac{1}{\sum_{\hat{\mathbf{x}}\in\hat{\mathcal{X}_{t}}}\hat{w}(\hat{\mathbf{x}})}p_{\hat{w}}(\hat{\mathcal{X}_{t}})\right),$ (13)
where $p_{\hat{w}}(\hat{\mathcal{X}_{t}})=\sum_{\hat{\mathbf{x}}\in\hat{\mathcal{X}_{t}}}\hat{w}(\hat{\mathbf{x}})\delta_{\hat{\mathbf{x}}}$, ${\rm\mathcal{W}_{1}}$ is the 1-Wasserstein distance with Euclidean norm.
Total loss and algorithm.
Note that the matching loss can be easily generalized to multi-time snapshots by employing Eq. (12) among two consecutive snapshots $p_{t}$ and $p_{t+1}$. Combined with distribution fitting loss in Eq. (13), we obtain the final loss
$\mathcal{L}(\theta,\omega)=\mathcal{L}_{\mathrm{VGFM}}(\theta,\omega)+\mathcal{L}_{\mathrm{OT}}(\theta,\omega)$ (14)
During training, we employ a parameter scheduling scheme in which we initially use $\mathcal{L}_{\rm VGFM}$ as a warm-up stage, before enabling both loss terms for joint training. This design provides a notable advantage: after the warm-up stage, $v_{\theta}$ and $g_{\omega}$ are well-initialized through conditional matching, which facilitates convergence when switching to $\mathcal{W}_{1}$-based training. The detailed algorithm is presented in Algorithm 1. For further details on the warm-up procedure and parameter scheduling, please refer to Appendix B.1.
Input: Observed data $\mathcal{X}_{0},\dots,\mathcal{X}_{T-1}$; warm-up epochs $M_{1}$; training epochs $M_{2}$
Output: Trained velocity field $v_{\theta}$ and growth function $g_{\omega}$
Compute transport plans $\pi^{t\to t+1}$ for $t=0,\dots,T-2$;
for $i=1$ to $M_{2}$ do
       for $t_{0}=0$ to $T-2$ do
             Sample a batch $(i,j)\sim\pi^{t_{0}\to t_{0}+1}$;
             Sample $t\sim\mathcal{U}(0,1)+t_{0}$, and $\mathbf{x}_{t}\sim\mathcal{N}(t\mathbf{x}_{1}^{j}+(1-t)\mathbf{x}_{0}^{i},\sigma^{2}I)$;
             $\mathcal{L}\leftarrow\mathcal{L}+\left\|v_{\theta}(\mathbf{x}_{t},t)-(\mathbf{x}_{1}^{j}-\mathbf{x}_{0}^{i})\right\|^{2}+\left|g_{\omega}(\mathbf{x}_{t},t)-\log([\pi^{0\to 1}\mathbf{1}_{m}]_{i})\right|^{2}$;
             if $i>M_{1}$ then
                   $\hat{\mathcal{X}}_{t_{0}+1}\leftarrow\phi_{v_{\theta}}(\hat{\mathcal{X}}_{t_{0}},t_{0}+1)$, $\hat{w}_{t_{0}+1}(\hat{\mathcal{X}}_{t_{0}+1})\leftarrow\phi_{g_{\omega}}(\hat{\mathcal{X}}_{t_{0}},t_{0}+1)$;
                   $\mathcal{L}\leftarrow\mathcal{L}+\mathrm{\mathcal{W}_{1}}\left(\frac{1}{N_{t_{0}+1}}p(\mathcal{X}_{t_{0}+1}),\frac{1}{\sum_{\hat{\mathbf{x}}\in\hat{\mathcal{X}_{t}}}\hat{w}(\hat{\mathbf{x}})}p_{\hat{w}}(\hat{\mathcal{X}}_{t_{0}+1})\right)$;
      Update $\theta,\omega$ using $\mathcal{L}$
Algorithm 1 Training algorithm of joint velocity-growth flow matching
4 Experiments
We evaluate VGFM on synthetic and real datasets, and compare it with state-of-the-art approaches, including methods that account for distributional unbalancedness [25, 49, 43] as well as those that assume mass conservation [21, 36, 39]. Our code will be released online.
Synthetic datasets. Inspired by [22, 25], we adopt the Simulation Gene dataset that applies a three-gene regulatory network to produce a quiescent region and an area exhibiting both transition and growth that can be observed. We also use Dyngen [52] to simulate a scRNA-seq dataset from a dynamic cellular process, which exhibits pronounced branching unbalancedness, with significantly different cell abundances across divergent lineages. Following the experiment setup of [30], we use PHATE [53] to reduce its dimensions to 5. Additionally, inspired by [25] that employs a high-dimensional Gaussian mixture model [54] to evaluate the scalability of models, we adopt a more challenging setting: 1000-dimensional Gaussian mixtures.
Real-world dataset. We conduct experiments on two real-world datasets, Embryoid Body (EB) [55] and CITE-seq (CITE) [56], both preprocessed following the procedures in [17, 21]. Each dataset is evaluated under both 5- and 50-dimensional PCA projections. Specifically, the EB (5D), CITE (5D), and CITE (50D) configurations are assessed using a hold-out strategy, same as [36], in which an intermediate time point is excluded during training. The model is then used to predict the distribution at the hold-out time, and we compute the $\mathcal{W}_{1}$ distance between the predicted and true distributions at that time point. The EB (50D) setting is further used for comparison against [25], evaluated using both $\mathcal{W}_{1}$ and the Relative Mass Error (RME) (see below).
Evaluation metrics. We follow [25] to use $\mathcal{W}_{1}$ to measure the distance between the predicted cell distribution and the true cell distribution. Additionaly, to assess the accuracy of the growth function $g_{\omega}$, we normalize the total mass at time 0 to 1, and denote the relative mass at time $t$ w.r.t time 0 as $m_{t}$, i.e., $m_{t}=\frac{N_{t}}{N_{0}}$. The predicted relative mass at time $t$ based on the evolution induced by $g_{\omega}$ is denoted as $\hat{m}_{t}=\sum_{\hat{\mathbf{x}}\in\hat{\mathcal{X}_{t}}}\hat{w}(\hat{\mathbf{x}})/N_{0}$, where $\sum_{\hat{\mathbf{x}}\in\hat{\mathcal{X}_{t}}}\hat{w}(\hat{\mathbf{x}})$ is the predicted total mass at time $t$. We define the relative mass error, denoted as RME, i.e., $\mathrm{RME}=\frac{|m_{t}-\hat{m}_{t}|}{m_{t}}$.
Inplementation details. We employ a 3-layer (5-layer for dimensions greater than 50) MLP with 256 hidden units and LeakyReLU activation to parameterize both the velocity field $v_{\theta}$ and the growth function $g_{\omega}$. Optimization is performed using the Adam optimizer with a learning rate of $10^{-4}$. A warm-up stage of 200-800 iterations is applied (only matching loss in Eq. (12) ), after which the distribution fitting loss $\mathcal{L}_{\mathrm{OT}}$ defined in Eq. (13) is applied for an additional 30 training epochs. We will discuss the strategy to set $\epsilon$ and $\tau$ in the Appendix B.1.
4.1 Results and Analysis
Ability to reconstruct cellular dynamics. As shown in Tab. 1, our model achieves lowest mean $\mathcal{W}_{1}$ and RME on synthetic datasets, outperforming both unbalancedness-aware methods [25, 22, 49] and unbalancedness-ignoring methods [36, 21], which demonstrate the ability of VGFM to model cellular dynamics for data and mass fitting. We also visualize the learned trajectory and growth function on the Simulation Gene dataset in Figs. LABEL:fig:6pic (a-c). Figures LABEL:fig:6pic (a-c) show that VGFM not only reconstructs accurate trajectories but also produces faithful predictions of spatially varying growth rates.
Scalability and mass-matching performance compared with existing unbalanced models. While existing SOTA approaches of UDSB [49], TIGON [22] and DeepRUOT [25] incorporate unbalanced modeling, these models either introduce biologically implausible transitions [49] (as discussed in [25]) or face difficulties in scaling to high-dimensions [49, 25, 22]. As a result, these methods do not perform well or converge on the 1000-dimensional Gaussian dataset, as in Tab. 1. In contrast, VGFM integrates flow matching based on unbalanced transport, allowing it to scale more effectively to high-dimensional datasets. Our method more accurately recovers the ground-truth dynamics defined in [25] and achieves superior mass-matching accuracy, as shown in Tab. 1 and Figs. LABEL:fig:6pic (d-f).
Table 1: Mean $\mathcal{W}_{1}$ and RME over all time on synthetic datasets. *OT-CFM and OT-MFM do not model the growth function, thus, RME is not computed. “N/C” means “not converge”.
Method Simulation Gene (2D) Dyngen (5D) Gaussian (1000D)
$\mathcal{W}_{1}$ RME $\mathcal{W}_{1}$ RME $\mathcal{W}_{1}$ RME
OT-CFM* [21] 0.302 3.926 10.126
OT-MFM* [36] 0.311 3.976 11.008
UDSB [49] 0.665 0.192 1.914 0.658 25.304 0.965
TIGON [22] 0.099 0.065 1.029 0.542 N/C N/C
DeepRUOT [25] 0.068 0.016 0.474 0.199 N/C N/C
VGFM 0.046 0.006 0.420 0.053 3.010 0.037
Hold-one-out results on real-world datasets. Table 2 compares hold-one-out results of different methods on EB (5D) and CITE (5D and 50D) datasets. It can be observed that VGFM outperforms most existing approaches in terms of $\mathcal{W}_{1}$ distance. We attribute this to the integration of complementary strengths from both flow matching and simulation-based (Neural ODE) frameworks. On the one hand, flow matching provides a robust initialization for learning the velocity field and growth rate, which helps maintain training stability after incorporating the distribution fitting loss $\mathcal{L}_{\mathrm{OT}}$, leading to better performance than purely simulation-based models, as also shown in Fig. LABEL:fig:eb_vism. On the other hand, since the flow matching objective serves only as an upper bound on the true Wasserstein distance [57], the introduction of $\mathcal{L}_{\mathrm{OT}}$ allows direct optimization of the distribution fitting loss, further enhancing model performance. Moreover, by explicitly modeling cell growth through a time-varying weight function, our framework generalizes beyond the conventional setting where Wasserstein distance is computed between unweighted discrete measures, thereby offering greater flexibility in evaluation and enabling more possible transport solutions.
Ablation study on loss terms. To assess the contribution of each loss component in our framework,
Table 2: Mean hold-one-out results on EB and CITE datasets over hold-out times.
Method EB CITE
5D 5D 50D
OT-CFM [21] 0.790 0.882 38.756
SF2M [39] 0.793 0.920 38.524
WLF-UOT [43] 0.738 0.733 37.007
UDSB [49] 1.206 2.023 44.168
OT-MFM [36] 0.713 0.724 36.394
DeepRUOT [25] 0.774 0.648 35.539
VGFM 0.676 0.650 34.182
we perform an ablation study on the EB dataset with unwhitened 50-dimensional PCA features. Table 3 reports the performance across four time points. VGFM in general achieves the lowest $\mathcal{W}_{1}$ and RME values (excluding time 1) compared with VGFM (w/o $\mathcal{L}_{\rm{OT}}$) and VGFM (w/o $\mathcal{L}_{\rm{VGFM}}$), demonstrating the effectiveness of both our joint velocity-growth matching loss and distribution fitting loss. Notably, by removing $\mathcal{L}_{\mathrm{VGFM}}$, both $\mathcal{W}_{1}$ and RME increase by a significant margin, implying the importance of our velocity and growth matching to our framework. Consistent with our motivation in Sect. 3.4, adding in distribution fitting loss $\mathcal{L}_{\mathrm{OT}}$ results in lower $\mathcal{W}_{1}$ and RME, improved fitting ability to observed distributions.
Comparison on computational efficiency. We compare the computational cost of our method and the SOTA method DeepRUOT [25] in Tab. 3. For fair comparison, all methods in Tab. 3 adopt the same networks 5-layer MLP with 256 hidden units, and are trained on the same GPU until convergence. As shown in Table 3, our full model requires slightly more training time than the variant using only $\mathcal{L}_{\rm VGFM}$, due to the training with additional distribution fitting loss. However, it still achieves faster convergence compared with simulation-based methods of DeepRUOT [25] and the ablation variant without $\mathcal{L}_{\rm VGFM}$. This demonstrates the critical role of the loss $\mathcal{L}_{\rm VGFM}$ for facilitating training convergence.
Table 3: Ablation study on EB (50D) dataset. For comparison, we also report the results of the SOTA approach DeepRUOT [25] in this table.
Method $t_{1}$ $t_{2}$ $t_{3}$ $t_{4}$ Training time
$\mathcal{W}_{1}$ RME $\mathcal{W}_{1}$ RME $\mathcal{W}_{1}$ RME $\mathcal{W}_{1}$ RME
DeepRUOT [25] 8.169 0.416 9.041 0.415 9.348 0.119 9.808 0.296 90 (mins)
VGFM (w/o $\mathcal{L}_{\rm{OT}}$) 8.915 0.020 10.590 0.098 10.915 0.067 11.635 0.088 6 (mins)
VGFM (w/o $\mathcal{L}_{\rm{VGFM}}$) 8.644 0.650 10.167 0.710 11.052 0.823 11.530 0.862 62 (mins)
VGFM 7.951 0.089 8.747 0.042 9.244 0.019 9.620 0.044 13 (mins)
5 Conclusion, Limitation, and Future Work
The paper proposes the joint Velocity-Growth Flow Matching (VGFM) method that jointly learns state transition and mass growth of single-cell populations via flow matching. VGFM designs an ideal dynamics containing a velocity field and a growth function, driven by our presented two-period dynamic understanding of the static semi-relaxed optimal transport model. Approximating the ideal dynamics using neural networks yields the velocity-growth joint flow matching framework. Extensive experiments on both synthetic and real datasets demonstrate the effectiveness of our method.
Although our approach achieves better scalability and training efficiency compared to simulation-based methods, it is not entirely simulation-free due to the incorporation of the distribution fitting loss $\mathcal{L}_{\rm OT}$. Moreover, since the learned growth rate relies on the observed cell counts at each time point, it may inadvertently capture effects that are not solely attributable to biological growth. Future work could address them by integrating biological priors into the learning process of the growth function $g$, and by exploring fully simulation-free alternatives that can achieve comparable performance.
Appendix C discusses the broader impacts of our work.
Appendix A Proofs
A.1 Background: Brenier-Benamou Formulation
We first recall the dynamic formulation of balanced optimal transport, which is known as Brenier-Benamou formula [58]. In this paper, we assume $p_{0}$ and $p_{1}$ are absolutely continuous w.r.t. Lebesgue measure, which is omitted for convenience of description.
Theorem A-2 (Brenier-Benamou formula).
Given two probability measures $p_{0},p_{1}\in\mathcal{P}_{2}(\Omega)$, it holds that
$\mathcal{W}_{2}^{2}(p_{0},p_{1})=\inf_{\rho_{t},v_{t}}\left\{\int_{0}^{1}\int_{\Omega}\|v_{t}(\mathbf{x})\|^{2}\mathrm{d}\mathbf{x}\mathrm{d}t\Big{|}\partial_{t}p_{t}=-\nabla\cdot(p_{t}v_{t}),\ p_{t=0}=p_{0},p_{t=1}=p_{1}\right\}$ (A-15)
and the optimal $v_{t}^{*}(\mathbf{x})$ can be expressed by Monge map between $p_{0}$ and $p_{1}$. i.e.,
$v_{t}(\mathbf{x}+t(T^{*}(\mathbf{x})-\mathbf{x}))=T^{*}(\mathbf{x})-\mathbf{x},\ \mathbf{x}\sim p_{0}$ (A-16)
A.2 Proof of Proposition 1
For the convenience of reading, here we restate Proposition 1 in the main text as Proposition A-2.
Proposition A-2.
Assume $c(\mathbf{x}_{0},\mathbf{x}_{1})=\|\mathbf{x}_{0}-\mathbf{x}_{1}\|^{2}$ and if we enforce ${\rm P}_{\#}^{0}\pi$ and $p_{0}$ to share the same support for admissible solution $\pi$ to problem (4), then we have $\min_{\pi}\mathcal{J}_{\rm sot}(\pi)=\min_{v_{t},g_{t}}\mathcal{J}^{\lambda}_{\rm tpt}(v_{t},g_{t}),\forall\lambda\in(0,1)$. Moreover, for any $\lambda\in(0,1)$, given the optimal transport plan $\pi^{*}$ to problem (4), let $p_{\lambda}^{*}\triangleq{\rm P}_{\#}^{0}\pi^{*}$, then $\pi^{*}$ can be expressed as $\pi^{*}=(\mathrm{Id},T^{*})_{\#}p^{*}_{\lambda}$ where $T^{*}$ is the Monge map between $p^{*}_{\lambda}$ and $p_{1}$. Meanwhile, there exist a $g_{t}^{*}$ such that $p_{\lambda}^{*}=p_{0}(\mathbf{x})e^{\int_{0}^{\lambda}g^{*}_{t}(\mathbf{x})\mathrm{d}t}$, and a $v_{t}^{*}$ given by
$v^{*}_{t}\left(\mathbf{x}+\frac{t-\lambda}{1-\lambda}(T^{*}(\mathbf{x})-\mathbf{x})\right)=\frac{T^{*}(\mathbf{x})-\mathbf{x}}{1-\lambda},$ (A-17)
satisfying $(v_{t}^{*},g_{t}^{*})\in\arg\min_{v_{t},g_{t}}\mathcal{J}^{\lambda}_{\rm tpt}(v_{t},g_{t})$.
Proof.
Recall that when $c(\mathbf{x}_{0},\mathbf{x}_{1})=\|\mathbf{x}_{0}-\mathbf{x}_{1}\|^{2}$, problems (4) and (5) are respectively
$\min_{\pi\geq 0}\mathcal{J}_{\rm sot}(\pi)\triangleq\int_{\Omega^{2}}\|\mathbf{x}_{0}-\mathbf{x}_{1}\|^{2}\ \,\mathrm{d}\pi(\mathbf{x}_{0},\mathbf{x}_{1})+\mathrm{KL}({\rm P}_{\#}^{0}\pi\|p_{0})\quad\text{subject to}\quad{\rm P}_{\#}^{1}\pi=p_{1},$
and
$\min_{(v_{t},g_{t})\in\mathcal{C}_{\lambda}(p_{0},p_{1})}\mathcal{J}^{\lambda}_{\rm tpt}(v_{t},g_{t})\triangleq(1-\lambda)\int_{\Omega}\int_{\lambda}^{1}p_{t}(\mathbf{x})\|v_{t}(\mathbf{x})\|^{2}\mathrm{d}t\mathrm{d}x+\mathcal{H}(v_{t},g_{t},p_{t}),$
where $\mathcal{H}(v_{t},g_{t},p_{t})=\int_{\Omega}p_{0}(\mathbf{x})(e^{\int_{0}^{\lambda}g_{t}(\mathbf{x})\mathrm{d}t}(\int_{0}^{\lambda}g_{t}(\mathbf{x})\mathrm{d}t-1)+1)\mathrm{d}x$, $\mathcal{C}_{\lambda}(p_{0},p_{1})=\{(v_{t},g_{t}):\partial_{t}p_{t}=g_{t}p_{t},t\in[0,\lambda];\partial_{t}p_{t}=-\nabla\cdot(p_{t}v_{t}),t\in(\lambda,1]\}$
(1) We first prove $\min_{\pi}\mathcal{J}_{\rm sot}(\pi)\geq\min_{v_{t},g_{t}}\mathcal{J}^{\lambda}_{\rm tpt}(v_{t},g_{t})$.
Let $\pi^{*}$ be the optimal solution to problem (4). For any $\lambda\in(0,1)$, define $p_{\lambda}={\rm P}^{0}_{\#}\pi^{*}$, and let $w_{\lambda}(\mathbf{x}):=\frac{p_{\lambda}(\mathbf{x})}{p_{0}(\mathbf{x})}$. Define the growth as
$\frac{\mathrm{d}}{\mathrm{d}t}\log w_{t}(\mathbf{x})=g_{t}(\mathbf{x}),t\in[0,\lambda].$
[…]
By Brenier’s theorem [51], we have $\pi^{*}=(\mathrm{Id},T^{*})_{\#}p_{\lambda}$, where $T^{*}(\mathbf{x})$ is the Monge map from $p_{\lambda}$ to $p_{1}$. Define
$g_{t}(\mathbf{x})=\frac{\log p_{\lambda}(\mathbf{x})-\log p_{0}(\mathbf{x})}{\lambda},\quad v_{t}(\mathbf{x}+\frac{t-\lambda}{1-\lambda}(T^{*}(\mathbf{x})-\mathbf{x}))=\frac{T^{*}(\mathbf{x})-\mathbf{x}}{1-\lambda}.$ (A-18)
Plugging into the dynamic objective and using the definition of KL-divergence:
$\mathrm{KL}(p_{\lambda}\|p_{0})=\int_{\Omega}p_{0}(\mathbf{x})\left(\frac{p_{\lambda}(\mathbf{x})}{p_{0}(\mathbf{x})}\left(\log\frac{p_{\lambda}(\mathbf{x})}{p_{0}(\mathbf{x})}-1\right)+1\right)\mathrm{d}\mathbf{x},$
we obtain
$\mathcal{J}^{\lambda}_{\rm tpt}(v_{t},g_{t})=$ $\int_{\Omega}p_{0}(\mathbf{x})\left(e^{\int_{0}^{\lambda}g_{t}(\mathbf{x})\mathrm{d}t}\left(\int_{0}^{\lambda}g_{t}(\mathbf{x})\mathrm{d}t-1\right)+1\right)\mathrm{d}\mathbf{x}+(1-\lambda)\int_{\Omega}\|T^{*}(\mathbf{x})-\mathbf{x}\|^{2}p_{\lambda}(\mathbf{x})\mathrm{d}\mathbf{x}$
$=$ $\mathrm{KL}(p_{\lambda}\|p_{0})+\int_{\Omega^{2}}\|\mathbf{x}_{0}-\mathbf{x}_{1}\|^{2}\mathrm{d}\pi^{*}(\mathbf{x}_{0},\mathbf{x}_{1})$
$=$ $\mathcal{J}_{\rm{sot}}(\pi^{*})$
Thus, $\mathcal{J}_{\rm sot}(\pi^{*})=\mathcal{J}^{\lambda}_{\rm tpt}(v_{t},g_{t})\geq\min_{v_{t},g_{t}}\mathcal{J}^{\lambda}_{\rm tpt}(v_{t},g_{t})$.
(2) To show the reverse inequality $\min_{\pi}\mathcal{J}_{\rm sot}(\pi)\leq\min_{v_{t},g_{t}}\mathcal{J}^{\lambda}_{\rm tpt}(v_{t},g_{t})$, we assume the contrary that $\min_{\pi}\mathcal{J}_{\rm sot}(\pi)>\min_{v_{t},g_{t}}\mathcal{J}^{\lambda}_{\rm tpt}(v_{t},g_{t})$. Let $(v_{t}^{*},g_{t}^{*})$ be an optimal solution to problem (5), and define
$\tilde{p}_{\lambda}(\mathbf{x}):=p_{0}(\mathbf{x})\exp\left(\int_{0}^{\lambda}g_{t}^{*}(\mathbf{x})\mathrm{d}t\right),$
Since the second-stage evolution $p_{t}$, for $t\in(\lambda,1]$, is governed solely by the velocity field $v_{t}$ and does not involve mass creation or destruction, both $\tilde{p}_{\lambda}$ and $p_{1}$ have the same total mass. Thus, the Monge map $\tilde{T}^{*}$ from $\tilde{p}_{\lambda}$ to $p_{1}$ under quadratic cost is well-defined. Then, by the Benamou–Brenier formulation [58, 59], we have
$v_{t}^{*}(\mathbf{x}+\frac{t-\lambda}{1-\lambda}(\tilde{T}^{*}(\mathbf{x})-\mathbf{x}))=\frac{\tilde{T}^{*}(\mathbf{x})-\mathbf{x}}{1-\lambda},$
and the corresponding coupling $\tilde{\pi}^{*}:=(\mathrm{Id},\tilde{T}^{*})_{\#}\tilde{p}_{\lambda}$ satisfies
$\mathcal{J}_{\rm sot}(\tilde{\pi}^{*})=$ $\mathrm{KL}(\tilde{p}_{\lambda}\|p_{0})+\int_{\Omega^{2}}\|\mathbf{x}_{0}-\mathbf{x}_{1}\|^{2}\mathrm{d}\tilde{\pi}^{*}(\mathbf{x}_{0},\mathbf{x}_{1})$
$=$ $\int_{\Omega}p_{0}(\mathbf{x})\left(e^{\int_{0}^{\lambda}g_{t}^{*}(\mathbf{x})\mathrm{d}t}\left(\int_{0}^{\lambda}g_{t}^{*}(\mathbf{x})\mathrm{d}t-1\right)+1\right)\mathrm{d}\mathbf{x}+(1-\lambda)\int_{\Omega}\|v_{t}^{*}(\mathbf{x})\|^{2}\tilde{p}_{\lambda}(\mathbf{x})\mathrm{d}\mathbf{x}$
$=$ $\mathcal{J}^{\lambda}_{\rm tpt}(v_{t}^{*},g_{t}^{*})=\min_{v_{t},g_{t}}\mathcal{J}^{\lambda}_{\rm tpt}(v_{t},g_{t})<\min_{\pi}\mathcal{J}_{\rm sot}(\pi)$
which leads to the contradiction.
Combining (1) and (2), we have $\mathcal{J}_{\rm sot}(\pi^{*})=\mathcal{J}^{\lambda}_{\rm tpt}(v_{t}^{*},g_{t}^{*})$.
Given $\pi^{*}$, we construct $g_{t}^{*},v_{t}^{*}$ as in Eq. (A-18). According to the proof of (1), we have $\mathcal{J}^{\lambda}_{\rm tpt}(v_{t}^{*},g_{t}^{*})=\mathcal{J}_{\rm sot}(\pi^{*})=\min_{v_{t},g_{t}}\mathcal{J}^{\lambda}_{\rm tpt}(v_{t},g_{t})$. ∎
A.3 Proof and Empirical Evidence of Theorem 1
For the convenience of reading, here we restate Theorem 1 in the main text as Theorem A-3.
Theorem A-3.
Gvien the initial distribution $p_{0}$, denote the ending distribution of the two-period dynamics
$\partial_{t}p_{t}=g_{t}p_{t},t\in[0,\lambda];\quad\partial_{t}p_{t}=-\nabla\cdot(p_{t}v_{t}),t\in(\lambda,1],$ (A-19)
as $p_{1}$, and denote the ending distribution of the joint dynamics starting from $p_{0}$
$\partial_{t}\tilde{p}_{t}=-\nabla\cdot(\tilde{p}_{t}\tilde{v}_{t})+\tilde{g}_{t}\tilde{p}_{t},\quad t\in[0,1],\quad\tilde{p}_{0}=p_{0},$ (A-20)
as $\tilde{p}_{1}$, then it holds that $\tilde{p}_{1}=p_{1}$.
Proof.
Given $\mathbf{x}_{0}\sim p_{0}$, consider the two systems below.
System I (original two-period transport dynamics):
$\begin{cases}\dfrac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}=v_{t}(\mathbf{x}_{t})\cdot\mathbb{I}_{(\lambda,1]}(t),\\ \dfrac{\mathrm{d}}{\mathrm{d}t}\log w_{t}(\mathbf{x}_{t})=g_{t}(\mathbf{x}_{t})\cdot\mathbb{I}_{[0,\lambda]}(t),\end{cases}$ (A-21)
where $w_{t}$ is the time-dependent weight function.
System II (joint dynamics defined via reparameterization (6)):
$\begin{cases}\dfrac{\mathrm{d}{\mathbf{x}}_{t}}{\mathrm{d}t}=\tilde{v}_{t}({\mathbf{x}}_{t}),\\ \dfrac{\mathrm{d}}{\mathrm{d}t}\log{w}_{t}({\mathbf{x}}_{t})=\tilde{g}_{t}({\mathbf{x}}_{t}),\end{cases}$ (A-22)
with $\tilde{w}_{t}$ being the time-dependent weight under joint dynamics. We first recall the definition $\tilde{v}_{t}(\mathbf{x})=(1-\lambda)\cdot v_{(1-\lambda)t+\lambda}(\mathbf{x}),\tilde{g}_{t}(\mathbf{x})=\lambda\cdot g_{\lambda t}\left(\psi_{\tilde{v},t}^{-1}(\mathbf{x})\right)$. To prove Theorem 1, it suffices to show that given the same initialization $\mathbf{x}_{0},w_{0}(\mathbf{x}_{0})$, the final state of system I ($\mathbf{x}_{1},w_{1}(\mathbf{x}_{1})$) and II ($\tilde{\mathbf{x}}_{1},\tilde{w}_{1}(\tilde{\mathbf{x}}_{1})$) are identical, i.e., $\tilde{\mathbf{x}}_{1}=\mathbf{x}_{1}$ and $\tilde{w}_{1}(\tilde{\mathbf{x}}_{1})=w_{1}(\mathbf{x}_{1})$.
$\tilde{\mathbf{x}}_{1}$ $=\mathbf{x}_{0}+\int_{0}^{1}\tilde{v}_{t}({\mathbf{x}})\,\mathrm{d}t$
$=\mathbf{x}_{0}+\int_{0}^{1}v_{(1-\lambda)t+\lambda}(\mathbf{x})\cdot(1-\lambda)\,\mathrm{d}t$
$=\mathbf{x}_{0}+\int_{\lambda}^{1}v_{s}\left(\mathbf{x}\right)\,\mathrm{d}s\quad(\mbox{let }s=(1-\lambda)t+\lambda)$
$=\mathbf{x}_{1}.$
Meanwhile,
$\log w_{1}(\mathbf{x}_{1})$ $=\log w_{\lambda}(\mathbf{x}_{\lambda})=\log w_{\lambda}(\mathbf{x}_{0})$
$=\log w_{0}(\mathbf{x}_{0})+\int_{0}^{\lambda}g_{s}(\mathbf{x}_{0})\,\mathrm{d}s$
$=\log w_{0}(\mathbf{x}_{0})+\int_{0}^{1}\lambda\cdot g_{\lambda t}(\mathbf{x}_{0})\,\mathrm{d}t\quad(\mbox{let }t=\frac{1}{\lambda}s)$
$=\log w_{0}(\mathbf{x}_{0})+\int_{0}^{1}\lambda\cdot g_{\lambda t}(\psi_{\tilde{v},t}^{-1}(\mathbf{x}))\,\mathrm{d}t\quad(\mathbf{x}=\mathbf{x}_{0}+\int_{0}^{t}\tilde{v}_{s}(\mathbf{x})\mathrm{d}s)$
$=\log w_{0}(\mathbf{x}_{0})+\int_{0}^{1}\tilde{g}_{t}(\mathbf{x})\,\mathrm{d}t$
$=\log\tilde{w}_{1}(\tilde{\mathbf{x}}_{1}).$
Hence, the final state and mass at $t=1$ under both systems coincide. Applying the above analysis to all $\mathbf{x}_{0}\sim p_{0}$ completes the proof. ∎
Empirical evidence of Theorem 1.
We provide the following example as an evidence to Theorem 1. Set $p_{0}=\mathcal{N}(2,0.5),\ v_{t}(\mathbf{x})=2t,\ g_{t}(\mathbf{x})=-\log(\mathbf{x}+1)+t^{3}$, using Eq.(6), we have $\tilde{v}_{t}(\mathbf{x})=(1-\lambda)v_{(1-\lambda)t+\lambda}(\mathbf{x}),\ \tilde{g}_{t}(\mathbf{x})=\lambda g_{\lambda t}(\psi_{\tilde{v},t}^{-1}({\mathbf{x}}))=\lambda g_{\lambda t}(\mathbf{x}-((1-\lambda)t+\lambda)^{2}+\lambda^{2})$. By setting $\lambda=0.4$ and using numerical solvers to solve the corresponding continuity equation, we validate our correctness of our theorem. As shown in Fig. 11, the two dynamics ended in the same distribution.
Figure A-11: Empirical evidence of Theorem 1.(a) original two-period transport dynamics. (b) joint dynamics defined via reparameterization 6. (c) the two dynamics ended in the same distribution.
A.4 Choosing $\tilde{g}_{t}$ According to Eq. (9) Minimizes the $L^{2}$-norm Energy Functional
We aim to prove the following statement: Given a function $w_{t}(\mathbf{x})$ whose logarithmic derivative satisfies $\frac{\mathrm{d}}{\mathrm{d}t}\log w_{t}(\mathbf{x})=g_{t}(\mathbf{x})$ and the boundary condition
$\log w_{1}(\mathbf{x})-\log w_{0}(\mathbf{x})=\int_{0}^{1}g_{t}(\mathbf{x})\,\mathrm{d}t,$
then among all functions $g_{t}(\mathbf{x})$ satisfying this constraint, the one minimizing the energy functional
$\mathcal{E}(g):=\int_{\Omega}\int_{0}^{1}\|g_{t}(\mathbf{x})\|_{2}^{2}\,\mathrm{d}t\mathrm{d}\mathbf{x}$
is the constant function $\int_{0}^{1}g_{t}(\mathbf{x})\mathrm{d}\mathbf{x}=\log w_{1}(\mathbf{x})-\log w_{0}(\mathbf{x}).$ To show this, we apply the method of Lagrange multipliers. Introduce a multiplier $\lambda(\mathbf{x})$ and consider the Lagrangian
$\mathcal{L}(g_{t})=\int_{\Omega}(\int_{0}^{1}\left(g_{t}(\mathbf{x})^{2}+\lambda(\mathbf{x})\,g_{t}(\mathbf{x})\right)\mathrm{d}t-\lambda(\mathbf{x})\left(\log w_{1}(\mathbf{x})-\log w_{0}(\mathbf{x})\right))\mathrm{d}\mathbf{x}.$
Taking the first variation of $\mathcal{L}$ with respect to $g_{t}(\mathbf{x})$ yields the optimality condition
$\frac{\delta\mathcal{L}}{\delta g_{t}}(\mathbf{x})=2g_{t}(\mathbf{x})+\lambda(\mathbf{x})=0\ \Rightarrow\ g_{t}(\mathbf{x})=-\frac{\lambda(\mathbf{x})}{2}.$
This shows that the optimal $g_{t}(\mathbf{x})$ is constant with respect to time. Plugging this into the constraint gives $g_{t}(\mathbf{x})=\log w_{1}(\mathbf{x})-\log w_{0}(\mathbf{x}).$ Hence, the constant function $g_{t}(\mathbf{x})$ is the unique minimizer of the energy functional under the given constraint.
Appendix B Experimental Details
The model architecture has been described in Sect. 4. We next detail the strategy for selecting the hyperparameters $\epsilon$ and $\tau$ in Eq. (11), as well as provide a comprehensive description of each experiment. All experiments are performed on a single-core CPU without GPU acceleration, and all visualizations are based on projections onto the first two dimensions of the high-dimensional data.
B.1 Determine the Entropy Regularization Parameter $\epsilon$ and Relaxation Coefficient $\tau$
Our algorithm involves solving a static semi-relaxed optimal transport problem (11), which includes two critical hyperparameters $\epsilon$ and $\tau$. These parameters play a pivotal role in the behavior and stability of the model. On one hand, if $\epsilon$ is too small, it may lead to numerical instability; if it is too large, it can cause incorrect cross-branch matching in unbalanced data scenarios. To select $\epsilon$, we first set $\tau$ to a moderately large value, such as $\tau=50$, and then perform a grid search over increasing values of $\epsilon$ starting from 0 until numerical stability is achieved. The choice of $\tau$ is even more crucial. When $\tau$ is too small, the KL divergence term between $P_{\#}^{0}\pi$ and $p_{0}$ receives a weak penalty. As a result, the transport mass becomes overly concentrated around points close to $p_{1}$, leading to highly uneven marginals $\pi\mathbf{1}_{m}$ in the discrete setting. This causes instability in the learning of the growth function $g$ and may result in erroneous modeling. Conversely, as $\tau\to+\infty$, Eq. (11) reduces to a balanced OT problem, which contradicts the unbalanced nature of our formulation. To determine a suitable value for $\tau$, we fix the previously selected small $\epsilon$ and gradually increase $\tau$ while observing the variation in $\sum\pi_{ij}c_{ij}$. This curve typically exhibits an increasing-then-flattening trend. Similar to the “elbow metho” used in clustering to select the optimal number of clusters, we identify the region where the curve becomes stable and choose $\tau$ accordingly, as illustrated in Fig. LABEL:fig:tau.
On the simulation gene dataset, we fix $\epsilon=0.001$ and evaluate the model with different values of $\tau\in\{5,10,15,20\}$, all of which lie within the identified stable region. The corresponding model performance metrics are summarized in Tab. A-4. We observe that the models trained with $\tau$ in this range are stable, validating the effectiveness of our hyperparameter selection strategy.
Table A-4: Fix $\epsilon=0.001$, change different $\tau$ on simulation gene data, we present the Wasserstein-1 distance of each timepoint.
Models $t_{1}$ $t_{2}$ $t_{3}$ $t_{4}$
$\mathcal{W}_{1}$ RME $\mathcal{W}_{1}$ RME $\mathcal{W}_{1}$ RME $\mathcal{W}_{1}$ RME
VFGM ($\tau=5$) 0.046 0.007 0.062 0.001 0.053 0.003 0.063 0.003
VFGM ($\tau=10$) 0.041 0.005 0.053 0.005 0.038 0.007 0.040 0.008
VFGM ($\tau=15$) 0.045 0.003 0.056 0.003 0.046 0.005 0.052 0.004
VFGM ($\tau=20$) 0.043 0.007 0.057 0.004 0.045 0.007 0.058 0.011
B.2 Simulation of Gene Data
Following the setup of [25], the dynamics of the simulated gene regulatory network are governed by the following system of differential equations:
$\frac{\mathrm{d}X_{1}}{\mathrm{d}t}$ $=\frac{\alpha_{1}X_{1}^{2}+\beta}{1+\alpha_{1}X_{1}^{2}+\gamma_{2}X_{2}^{2}+\gamma_{3}X_{3}^{2}+\beta}-\delta_{1}X_{1}+\eta_{1}\xi_{t},$ (A-23)
$\frac{\mathrm{d}X_{2}}{\mathrm{d}t}$ $=\frac{\alpha_{2}X_{2}^{2}+\beta}{1+\gamma_{1}X_{1}^{2}+\alpha_{2}X_{2}^{2}+\gamma_{3}X_{3}^{2}+\beta}-\delta_{2}X_{2}+\eta_{2}\xi_{t},$
$\frac{\mathrm{d}X_{3}}{\mathrm{d}t}$ $=\frac{\alpha_{3}X_{3}^{2}}{1+\alpha_{3}X_{3}^{2}}-\delta_{3}X_{3}+\eta_{3}\xi_{t},$
where $X_{i}(t)$ denotes the expression level of gene $i$ at time $t$. Genes $X_{1}$ and $X_{2}$ mutually inhibit each other while self-activating, forming a toggle-switch regulatory motif. An external signal $\beta$ independently activates both $X_{1}$ and $X_{2}$, whereas $X_{3}$ inhibits the expression of both $X_{1}$ and $X_{2}$.
The parameters $\alpha_{i}$ and $\gamma_{i}$ determine the strengths of self-activation and inhibition, respectively. The $\delta_{i}$ terms denote degradation rates, and $\eta_{i}\xi_{t}$ represents additive Gaussian noise with intensity $\eta_{i}$. The cell division rate is positively correlated with $X_{2}$ expression and is calculated as
$g=\frac{\alpha_{2}X_{2}^{2}}{1+X_{2}^{2}}\%.$
At each cell division event, daughter cells inherit gene expression values $(X_{1},X_{2},X_{3})$ from the parent, subject to small perturbations $\eta_{d}\mathcal{N}(0,1)$ per gene. Post-division, cells evolve independently according to the same stochastic dynamics.
Initial Conditions and Simulation Setup.
Initial expression states are sampled from two normal distributions: $\mathcal{N}([2,0.2,0],0.01)$ and $\mathcal{N}([0,0,2],0.01)$. At every step, negative values are clipped to zero. Gene expression data is recorded at time points $t\in\{0,8,16,24,32\}$.
By setting $\tau=10$ and $\epsilon=0.001$ according to our selection scheme described in Appendix B.1 and illustrated in Fig. LABEL:fig:tau (a), we train the model with a warm-up stage of 200 iterations. After that, distribution fitting loss $\mathcal{L}_{\rm OT}$ is introduced for an additional 30 training epochs. Our model achieve the best performance in terms of (weighted) $\mathcal{W}_{1}$ and relative mass error as shown in Tab. 1.
Table A-5: Simulation parameters on gene regulatory network [25].
Parameter Value Description
$\alpha_{1}$ 0.5 Strength of self-activation for $X_{1}$
$\gamma_{1}$ 0.5 Inhibition of $X_{2}$ by $X_{1}$
$\alpha_{2}$ 1 Strength of self-activation for $X_{2}$
$\gamma_{2}$ 1 Inhibition of $X_{1}$ by $X_{2}$
$\alpha_{3}$ 1 Strength of self-activation for $X_{3}$
$\gamma_{3}$ 10 Half-saturation constant for inhibition
$\delta_{1}$ 0.4 Degradation rate for $X_{1}$
$\delta_{2}$ 0.4 Degradation rate for $X_{2}$
$\delta_{3}$ 0.4 Degradation rate for $X_{3}$
$\eta_{1}$ 0.05 Noise intensity for $X_{1}$
$\eta_{2}$ 0.05 Noise intensity for $X_{2}$
$\eta_{3}$ 0.05 Noise intensity for $X_{3}$
$\eta_{d}$ 0.014 Noise for perturbation during cell division
$\beta$ 1 External activation signal
$\mathrm{d}t$ 1 Simulation time step
Time Points $\{0,8,16,24,32\}$ Observation time points
B.3 Dyngen Data
Dyngen is a multi-modal simulation engine for studying dynamic cellular processes at single-cell resolution [52]. We follow the setup of [30], which simulates gene expression time-series data that mimics cell proliferation processes, including branching and temporal progression. In the analyzed instance of the dyngen dataset, the samples span five discrete time points, labeled from 0 to 4, capturing the temporal evolution of the cell population.
We use PHATE [53] to reduce its dimensions to 5, the same as [30]. Starting from time point 0, cells can be divided into two branches based on the sign of the second PHATE coordinate ($x_{2}$). Quantitative analysis reveals a pronounced branch imbalance: for instance, at time point 4, 88 cells belong to one branch while 213 belong to the other, which poses a great challenge to the models that do not consider unbalancedness, resulting in cross-branch inference. Throughout this study, we assume that the temporal change in cell counts within each branch is entirely governed by a growth function, without contributions from migration or observational noise.
By setting $\tau=5$ and $\epsilon=0.03$ according to our selection scheme described in Appendix B.1 and illustrated in Fig. LABEL:fig:tau (b), we train the model with a warm-up stage of 400 iterations. After that, distribution fitting loss $\mathcal{L}_{\rm OT}$ is introduced for an additional 30 training epochs. Our model successfully avoids cross-branch reconstructions in trajectory modeling, as illustrated in Fig. LABEL:fig:dygen_vis. This leads to generated samples that remain well-aligned with the underlying manifold. In terms of growth rate estimation, our model also accurately captures the rapid expansion observed in the branch below between time point 3 and time point 4, where the number of observed cells increases dramatically from 25 to 213, indicating the highest growth rate in the system.
B.4 Gausssian 1000D Data
DeepRUOT [25] employs a high-dimensional Gaussian mixture distribution (100D) [54] to evaluate the scalability of models. In this work, we adopt an even more challenging setting by testing scalability at 1000 dimensions. Following the setup of [25], for the initial distribution, we generated 400 samples from the Gaussian located lower in the $(x_{1},x_{2})$ plane, and 100 samples from the Gaussian positioned higher. For the final distribution, we generated 1,000 samples from the upper Gaussian, and 200 samples each from the two lower Gaussians, and assume cells in the upper region exhibit proliferation without transport [5]. We observe that simulation-based methods [25] encounter training instability in the 1000-dimensional setting and flow matching-based methods fail to learn a reliable velocity field due to the unbalancedness of data, as shown in Tab. 1. (Some works incorporating unbalancedness [35, 40] may as well reconstruct correct velocity field but fail to construct growth rate). In contrast, By setting $\tau=5$ and $\epsilon=0.03$ according to our selection scheme described in Appendix B.1 and illustrated in Fig. LABEL:fig:tau (c), we train the model with a warm-up stage of 500 iterations. After that, distribution fitting loss $\mathcal{L}_{\rm OT}$ is introduced for an additional 10 training epochs. Our method remains effective even in high-dimensional regimes for both velocity and growth field learning.
B.5 Additional Experiment: Mouse Hematopoiesis Data
We also validate our algorithm on the mouse hematopoiesis data previously analyzed in [60, 22, 25]. This dataset leverages lineage tracing to track differentiation trajectories. After applying batch correction to integrate data across multiple experiments, the cells were embedded into a two-dimensional force-directed layout (SPRING plot). The resulting visualization reveals a pronounced bifurcation, where early progenitor cells diverge into two distinct differentiation lineages.
Our model effectively captures the underlying branching structure in the data, and the predicted growth rates align well with those reported in [25], as well as with established biological priors and are consistent with known biological lineage patterns. Following the evaluation protocol of [25], we also compute the Wasserstein-1 distance as a quantitative metric. Under this evaluation, by setting $\tau=20$ and $\epsilon=0.005$ according to our selection scheme described in Appendix B.1, we train the model with a warm-up stage of 200 iterations. After that, the distribution fitting loss $\mathcal{L}_{\rm OT}$ is introduced for an additional 50 training epochs. Our method achieves superior performance, demonstrating improved trajectory inference. In addition, we evaluate the RME (Relative Mass Error) metric introduced in this paper. The results indicate that our model significantly outperforms the other baselines in terms of mass-matching reconstruction accuracy, as shown in Tab. A-6 and Fig. A-16.
Table A-6: $\mathcal{W}_{1}$ and RME over all time on mouse hematopoiesis data. Part of the results were adopted from [25].
Models $t_{1}$ $t_{2}$
$\mathcal{W}_{1}$ RME $\mathcal{W}_{1}$ RME
SF2[39] 0.167 0.190
uAM[42] 0.745 0.777
UDSB [49] 0.388 0.159 0.128 0.249
TIGON [22] 0.314 0.124 0.342 0.177
DeepRUOT [25] 0.145 0.140 0.132 0.202
VGFM 0.115 0.043 0.094 0.019
Figure A-16: Comparison of predicted relative mass (UDSB, TIGON, DeepRUOT, VFGM) with observed values from mouse hematopoiesis data.
B.6 EB Data
Extensive experiments on the Embryoid Body (EB) [55] dataset with varying PCA dimensions are conducted in Sect. 4. We preprocess this dataset first by normalizing EB (5D) and leaving EB (50D) unnormalized. The proposed VFGM demonstrates strong performance in both cellular dynamics reconstruction and growth prediction, as illustrated in Fig. LABEL:fig:eb_vis. Specifically, we set the parameters as $\epsilon=0.01$, $\tau=30$, $M_{1}=800$, and $M_{2}=830$, and normalize the cost matrix to ensure numerical stability when computing the semi-relaxed optimal transport plans (Eq. (11)).
Big batch strategy for large-scale datasets. When the number of observed data points is large, computing the transport plan between two consecutive time points using Eq. (11) becomes computationally inefficient. To address this, we adopt a Big-Batch Strategy. Specifically, we partition the data at each time point into $n$ subsets, referred to as Big Batches. For each of the $n$ Big Batches, we precompute and store the transport plans between adjacent time points. During training, we randomly select one Big Batch and sample smaller mini-batches from it based on the precomputed transport plan for model optimization. In real-world datasets, we set $n=5$, while in synthetic datasets above we do not apply this strategy, i.e., $n=1$, as the sample size is relatively smaller. The mini-batch size is set to 256 for all experiments, except for Dyngen, where we use a smaller batch size of 32 due to its limited number of samples.
We also examine whether the distribution fitting loss might lead to overfitting on the given population snapshots. In the hold-one-out experiments (Tab. 2), we relabel the timestamps of the four timepoints as 0, 1, 2, 3, and 4. Following the protocol in [36], we evaluate the model by holding out one intermediate timepoint at a time (i.e., 1, 2, and 3) and report the average Wasserstein-1 distance. While hold-one-out experiments have already validated the superiority of VFGM over other methods, we further analyze the loss curves of both the distribution fitting loss and the $\mathcal{W}_{1}$ distance on a hold-out distribution (e.g., the distribution at time 1) that is not seen during training (Fig. A-18). The results show that the $\mathcal{W}_{1}$ distance on the hold-out distribution decreases during the optimization of $\mathcal{L}_{\rm OT}$, indicating the absence of overfitting and demonstrating the generalization ability of the distribution fitting loss in modeling unseen distributions.
Figure A-18: Loss curves of the training loss after incorporating $\mathcal{L}_{\rm OT}$ (blue) and the $\mathcal{W}_{1}$ distance on a hold-out distribution at timepoint 1 (red) on 5D EB data.
B.7 CITE Data
CITE-seq (Cellular Indexing of Transcriptomes and Epitopes by Sequencing) [56] is an advanced technique that enables the simultaneous profiling of transcriptomes and surface protein expression at the single-cell level through the use of antibody-derived tags. In this study, we utilize only the gene expression matrix from the CITE-seq dataset, and preprocess the data by normalizing CITE (5D) and leaving CITE (50D) unnormalized. The experimental settings for the CITE dataset are consistent with those used in the EB data experiments, employing PCA with 5 and 50 dimensions, and hyperparameters set as $\epsilon=0.01$, $\tau=30$, $M_{1}=800$, and $M_{2}=830$. We relabel the timestamps of the four timepoints as 0, 1, 2, and 3. Following the protocol in [36], we evaluate the model by holding out one intermediate timepoint at a time (i.e., 1 and 2) and report the average Wasserstein-1 distance.
To assess the model’s capability to capture state trajectories (Fig. LABEL:fig:cite_vis (a)) and predict mass growth (Fig. LABEL:fig:cite_vis (b)), the distribution at time point 1 is held out. Additionally, the distribution at time point 2 is held out to evaluate the performance of the $\mathcal{L}_{\rm OT}$ and $\mathcal{W}_{1}$ distance (hold-out) curves (Fig. LABEL:fig:cite_hoo). The strong performance observed in the visualizations substantiates the potential of the proposed VGFM framework.
Appendix C Broader Impacts
Our method provides a scalable and efficient framework for modeling cellular dynamics, enabling trajectory reconstruction and growth rate inference in high-dimensional single-cell datasets. In biomedical and clinical settings, such capabilities can facilitate a deeper understanding of developmental processes, disease progression, and response to treatment at a single-cell resolution. For example, modeling differentiation trajectories and proliferation patterns of stem or immune cells could inform therapeutic strategies in cancer, regenerative medicine, and immunotherapy.
However, the proposed method is inherently data-driven and relies on statistical patterns learned from observational data. As such, it may produce biologically implausible trajectories or growth behaviors that conflict with known biological priors, especially when the training data is noisy, biased, or incomplete. This could potentially lead to misleading interpretations or incorrect clinical hypotheses if not carefully validated by domain experts. Therefore, we emphasize that any downstream medical conclusions drawn from the model’s output should be interpreted with caution and in conjunction with biological prior knowledge and experimental validation.